OpenStudy (anonymous):

Which can not be written as an equivalent improper fraction? fourteen-fourths two and two-thirds one-sixth nine-eighths

OpenStudy (anonymous):

?

OpenStudy (anonymous):

im confuse can you help me ?

OpenStudy (anonymous):

In improper fractions, the numerator is larger than the denominator. Let's check each one: \[\frac{ 14 }{ 4 }\]^ Improper fraction\[2\frac{ 2 }{ 3 } = \frac{ 8 }{ 3 } \]^Can be changed into an improper fraction\[\frac{ 1 }{ 6 }\]^NOT an improper fraction.\[\frac{ 9 }{ 8 } \]^Improper fraction

OpenStudy (anonymous):

@melody I hope this helps! (: